# Asymptotic-preserving semi-implicit finite volume scheme for Extended Magnetohydrodynamics

### v1

@@ Abstract
A Finite Volume (FV) scheme is developed for solving the extended magnetohydrodynamic (XMHD) equations, yielding accurate results in the ideal, resistive, and Hall MHD limits. This is accomplished by first re-writing the XMHD equations such that it allows the algorithm to retain the use of ideal MHD Riemann solvers and the constrained transport method to preserve divergence-free magnetic fields. Incorporation of electron inertia and displacement current introduces additional numerical stiffness which motivates a semi-implicit FV scheme that re-formulates the XMHD model as a relaxation system. The equations are then advanced in time using an explicit 2nd-order Runge–Kutta scheme with operator splitting applied to the implicit source term updates at each sub-stage. For additional numerical stability, density-dependent slope limiter is implemented to increase flux diffusivity at low density where non-ideal effects become significant. The algorithm is subsequently implemented on Artemis, a multifluid radiation hydrodynamics code built on Parthenon framework for high-performance computing (HPC) and adaptive mesh refinement (AMR) support. As the new algorithm retains many aspects of the ideal MHD formulations, it asymptotes naturally to the ideal MHD limit. Moreover, it shows promising results at the resistive and Hall MHD limits. This is verified against published test problems for ideal, resistive and Hall MHD.
@@
Extended Magnetohydrodynamics , Field diffusion , Hall drift waves , Resistive MHD
@@ 2.1 Reformulation of governing equations
The above derivation can be found in A.
@@
$\frac{\partial}{\partial t}(\rho\mathbf{u})+\nabla\cdot\left[\rho\mathbf{u}\mathbf{u}-\mathbf{B}\mathbf{B}+\left(P+\frac{1}{2}B^{2}\right)\mathbb{I}\right]$ $=\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}\left(\mathbf{J}-\nabla\times\mathbf{B}\right)\times\mathbf{B}$ (18)
$\frac{\partial}{\partial t}\left(E_{n}+\frac{1}{2}B^{2}\right)+\nabla\cdot\left[\mathbf{u}\left(E_{n}+P+B^{2}\right)-(\mathbf{B}\cdot\mathbf{u})\mathbf{B}\right]$ $=\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}\left(-\mathbf{B}\cdot[\nabla\times(\mathbf{E}+\mathbf{u}\times\mathbf{B})]\right$.
$\left.+\mathbf{u}\cdot\left(\left(\mathbf{J}-\nabla\times\mathbf{B}\right)\times\mathbf{B}\right)+\eta J^{2}\right)$ (19)
@@
where $\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}$ is the coefficient for the explicit source terms that arises from non-dimensionalization and the dimensional constants should be chosen such that $\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}$ is $\sim$1. This does not necessary yield the best result which will be shown in one of the test case in Section 4.2 below but it serves as a good starting point to simplify the complexity of the above equations.
@@ 2.2 Strongly collisional plasma
One limitation of the above set of XMHD equations is that it does not contain source term to equilibrate ions and electrons temperature as the plasma becomes strongly collisional. To account for this, a more general equation for the electron entropy equation is considered
@@
The quasineutrality assumption then allows the electron-ion interaction to be written as [18]
@@
where the electron-ion exchange coefficient rate is $\nu_{ei}^{g}=\frac{3}{2}k_{B}\frac{n_{e}n_{i}}{\tau_{ie}n_{e}+\tau_{ei}n_{i}}$ and take $\nu_{ie}=\frac{n_{e}m_{e}}{n_{i}m_{i}}\nu_{ei}$ by assuming that both ions and electrons are close to thermal (Maxwellian) distributions where their momentum exhanges are symmetric. As seen from Eq. 7, Zhao et. al. [16] drops the $\nu_{ei}^{g}$ term and takes $\xi_{ei}$ as 1 by assuming that Ohmic heating between electrons and ions is similar. Therefore, some of the problem parameters used in the shocktube test case are chosen by Zhao et. al. such that the electron-ion energy exchange source term is weak. As a result, the exchange term is neglible over the simulation time, and $T_{i}$ and $T_{e}$ are not expected to equilibrate. We use the same parameter regime when testing the XMHD algorithm at the ideal MHD limit in this paper.
@@ 3.1 Slope limiter
where $\left(\frac{\lambda_{0}}{\Delta x}\right)^{3}$ is a free parameter and from the shock tube problem, 0.005 is a suitable value. This can be seen akin to a local Knudsen number of around 0.17 which lies in the transition regime away from the continuum limit. Therefore, as the flow density decreases and approaches the molecular limit, the above limiter approaches zero, giving rise to a 1st-order diffusive flux to stabilize the solution by damping out small amplitude but high frequency oscillations. Future work could be focused on a more rigorous density-dependent slope limiter without any free parameter.
@@ 3.2 Primitive form for flux Jacobian
Note that the explicit source terms depend on $\mathbf{E}$ and $\mathbf{J}$ which are implicitly updated by the diffusive backward Euler to relax the stiff source term to the equilibrium value where the characteristic speeds of the flux Jacobian becomes more important. Moreover, the relaxation accounts for the speed of light according to Eqs. 51 which shows that larger $\Delta t$ results in faster relaxation since more higher-order terms are damped out. However, these are not sufficient and thus, density-dependent slope limiters are required for the shock tube problem to further damp out small and fast oscillations. More rigorous analysis and testing will be needed in future.
@@ 4.2 Ideal MHD linear wave test
As expected, the results in Fig. 6(a) show that without Hall and resistive terms, the coupling of ideal MHD and other components through the implicit source term update is correct in ensuring that $\mathbf{E}\approx-\mathbf{u}\times\mathbf{B}$ due to the 2nd-order convergence from employing modified PLM and RK2 time integration. However, the entropy wave ceases to exhibit second-order convergence at very fine resolutions. This may result from the solver not sufficiently relaxing towards the ideal MHD limit as the grid resolution increases. Consequently, $\mathbf{J}$ does not achieve second-order convergence towards $\nabla\times\mathbf{B}$ at higher resolutions, causing the momentum source term to fail to converge to zero and thereby contaminating the entropy wave. For instance, when a different $L_{0}$ is used such that $\epsilon_{J}$ is smaller, as shown in Table 2, $\mathbf{J}$ converges more rapidly towards $\nabla\times\mathbf{B}$. This behavior is evident in Fig. 6(b), where using $L_{0}=10^{-2}$ leads to second-order convergence, unlike the original case with $L_{0}=10^{-3}$ at high resolution.
However, increasing $L_{0}$ to reduce $\epsilon_{J}$ is not ideal either, as it simultaneously increases $\epsilon_{E}$, making it more difficult for the displacement current to approach zero. This can also be observed in Fig. 6(b), where $L_{0}=10^{-1}$ yields the smallest $\epsilon_{J}$ but the largest $\epsilon_{E}$, thus generally have the largest RMS error, particularly at low resolutions. Therefore, an optimal choice of $L_{0}$ is required to ensure that both $\epsilon_{E}$ and $\epsilon_{J}$ remain sufficiently small for the problem at hand. This balance will be investigated further in future work. As mentioned earlier, while non-dimensionalizing the equations, ensuring that $\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}\approx 1$ does not always yield the best results (e.g., $L_{0}=10^{-3}$ vs $L_{0}=10^{-2}$ here), but this scaling will be adopted throughout the remainder of the paper for simplicity and consistency with previous studies.
$L_{0}=10^{-3}$ $L_{0}=10^{-2}$ $L_{0}=10^{-1}$
$\epsilon_{E}$ $1\times 10^{-9} 1.11×1$$\times 10^{-7} 1.1$$11\times 10^{-5}$
$\epsilon_{J}$ $4.71\times 10^{-10}$ $4.71\times 10^{-12}$ $4.71\times 10^{-14}$
$\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}$ $95\times 10^{-1} 9.9$$5\times 10^{-3} 9.9$$95\times 10^{-5}$
Table 2: Different relaxation parameters and explicit source term coefficient at different $L_{0}$.
@@
(b) Since entropy wave did not fully converge to 2nd-order especially at high resolution, a different $L_{0}$ is used to ensure that the XMHD solver is able to relax sufficiently to the ideal MHD limit for 2nd-order convergence. Note that when different $L_{0}$ is used, the background flow velocity and computational domain must be adjusted such that they are all consistent dimensionally.
@@ 4.3 Field Diffusion
where $x_{0}=10$cm, $\Delta x=3$cm and Atwood number is set as $A=20$ with a low number density of $n_{0}=10^{18}\text{m}^{-3}$ where Hall effect is significant. This means that the dimensional constants for the PDEs are also adjusted such that $n_{0}=10^{18}$, $L_{0}=1$ and $B_{0}$ set to 0.4568 such that the dimensionless coefficient ($\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}$) for the explicit source terms is approximately 1. The density profile is then maintained by setting $u_{x}=0$ every timestep. The plasma is also assumed to be isothermal ($T\equivconst$.) and for the ideal MHD components, HLLD Riemann solver [23] is used. Magnetic field is set along the z direction where $B_{0}=0.1\text{T}$ with perturbation of $\delta B_{z}=\delta B\cos\left(\frac{2\pi my}{L_{y}}\right)$ in the y-direction where $\delta B=0.001\text{T}$ and m is the mode number. The plasma beta ($\beta$) is then set to 0.0001 at peak density region where $x=0$. This allows the Hall drift wave to propagate in the y-direction.
@@
Rotor problem is also included to study the onset and propagation of strong torsional Alfven waves in both ideal and non-ideal regimes. Parthenon’s in-built adaptive mesh refinement (AMR), featuring a uniquely designed second-order prolongation and restriction scheme that preserves $\nabla\cdot\mathbf{B}\approx 0$ [36], is also tested with the XMHD solver. This is done using a computational domain of $0.5,0.5]\times[-0.5,0.5]$ with base grid resolution of 100 by 100 and 3 levels of adaptive grid refinement. The heat capacity ratio is set to $1.4$ here. For the ideal MHD case, $\rho_{0}$ is 1 and to test for the increment of non-ideal effects, $\rho_{0}$ is decreased to 0.1.
@@ 5 Conclusion
The development of a semi-implicit finite-volume schemes for solving the XMHD equations while retaining most of the already well-established ideal MHD framework has been the subject of the work. The goal is to ensure that the XMHD solver is asymptotically preserving in terms of producing similar results to existing ideal MHD solver. Additionally, the XMHD solver developed here has shown promising results at the resistive and Hall limits. This potentially enables the current solver to handle a wider range of density regimes, which is particularly important for high-energy-density (HED) plasma systems. Such systems often exhibit large density variations that must be accurately captured to improve one’s understanding and control of the plasma’s energy and stability.
@@
This is accomplished by first re-writing the governing equations such that it preserves most of the ideal MHD framework, allowing existing ideal MHD solvers or algorithms to be re-used. For instance, the advantage of preserving locally divergence-free magnetic field by placing magnetic-field at face-center through constrained transport method is employed in this XMHD solver. Naturally, the $\mathbf{E}$ field is placed at cell edges, together with $\mathbf{J}$ field for consistency during implicit source terms update. However, the implicit source term update requires directional splitting to avoid spatial biasing since it requires all of the edge-centered $\mathbf{J}$ and $\mathbf{E}$ components which are not co-located. The Jacobian matrix for the governing equations without the source terms is also obtained to determine the diffusive stabilization for the non-ideal components of the LLF Riemann solver evaluated at the cell faces. But this is not sufficient stability-wise, especially at low density where non-ideal effects become significant and therefore, a density-dependent slope limiter is used to increase flux diffusivity away from the continuum limit. Finally, since all the variables are not co-located, interpolation of values across different topological elements of the computational cell is required and simple arithmetic averaging seems sufficient for the test problems here.
@@
There are other aspects of the current algorithm that require further understanding and improvement. For example, the solver includes two stiff terms (i.e., $\epsilon_{E}$ and $\epsilon_{J}$) that depend on the dimensional constants used to non-dimensionalize the XMHD equations. At present, these constants are chosen such that $\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}\approx 1$, as this choice simplifies the problem; however, it has been shown that this does not necessarily yield the most accurate results. Instead, the ratio of the timestep $dt$ to $\epsilon_{E}$ and $\epsilon_{J}$ may influence how effectively the system relaxes towards the appropriate physical limit. Future work will involve additional tests and analysis to better understand this relationship. Furthermore, there are several potential avenues for enhancing the current algorithm such as

### v2

@@ Abstract
A Finite Volume (FV) scheme is developed for solving the extended magnetohydrodynamic (XMHD) equations, yielding accurate results in the ideal, resistive, and Hall MHD limits. This is accomplished by first re-writing the XMHD equations such that it allows the algorithm to retain the use of ideal MHD Riemann solvers and the constrained transport method to preserve divergence-free magnetic fields. Incorporation of electron inertia and displacement current introduces additional numerical stiffness which motivates a semi-implicit FV scheme that re-formulates the XMHD model as a relaxation system. The equations are then advanced in time using an explicit 2nd-order Runge–Kutta scheme with operator splitting applied to the implicit source term updates at each sub-stage. For additional numerical stability, a density-dependent slope limiter is implemented to increase flux diffusivity at low density regions where non-ideal effects become significant. The algorithm is subsequently implemented in a scalable adaptive mesh refinement (AMR) framework. As the new algorithm retains many aspects of the ideal MHD formulations, it asymptotes naturally to the ideal MHD limit. Moreover, it shows promising results at the resistive and Hall MHD limits. This is verified against reference test problems for ideal, resistive and Hall MHD.
@@
Extended Magnetohydrodynamics , Field diffusion , Hall drift waves
@@ 2.1 Reformulation of governing equations
The above derivation can be found in A. Extended comments on treatment of the electron entropy equation are provided in section 2.2.
@@
$\frac{\partial}{\partial t}(\rho\mathbf{u})+\nabla\cdot\left[\rho\mathbf{u}\mathbf{u}-\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}\mathbf{B}\mathbf{B}+\left(P+\frac{1}{2}\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}B^{2}\right)\mathbb{I}\right]$ (18)
$=\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}\left(\mathbf{J}-\nabla\times\mathbf{B}\right)\times\mathbf{B}$
$\frac{\partial}{\partial t}\left(E_{n}+\frac{1}{2}\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}B^{2}\right)+\nabla\cdot\left[\mathbf{u}\left(E_{n}+P+\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}B^{2}\right)-\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}(\mathbf{B}\cdot\mathbf{u})\mathbf{B}\right]$ (19)
$=\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}\left(-\mathbf{B}\cdot[\nabla\times(\mathbf{E}+\mathbf{u}\times\mathbf{B})]\right.\left.+\mathbf{u}\cdot\left(\left(\mathbf{J}-\nabla\times\mathbf{B}\right)\times\mathbf{B}\right)+\eta J^{2}\right)$
@@
where $\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}$ arises from non-dimensionalization and the dimensional constants should be chosen such that $\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}$ is $\sim$1. This ensures that the left-hand side of Eqs. 17 to 20 is equivalent to that of ideal MHD solvers.
@@ 2.2 Strongly collisional plasma
One limitation of the above set of XMHD equations is that it does not contain the source term responsible for equilibrating ion and electron temperatures as the plasma becomes strongly collisional. To account for this, a more general equation for the electron entropy equation is considered
@@
The quasi-neutrality assumption then allows the electron-ion interaction to be written as [18]
@@
where the electron-ion exchange coefficient rate is $\nu_{ei}^{g}=\frac{3}{2}k_{B}\frac{n_{e}n_{i}}{\tau_{ie}n_{e}+\tau_{ei}n_{i}}$ and take $\nu_{ie}=\frac{n_{e}m_{e}}{n_{i}m_{i}}\nu_{ei}$ by assuming that both ions and electrons are close to thermal (Maxwellian) distributions where their momentum exhanges are symmetric. As seen from Eq. 7, Zhao et. al. [16] drop the $\nu_{ei}^{g}$ term and take $\xi_{ei}$ as 1 by assuming that Ohmic heating between electrons and ions is similar. Therefore, some of the problem parameters used in the shocktube test case are chosen by Zhao et. al. such that the electron-ion energy exchange source term is weak. As a result, the exchange term is neglible over the simulation time, and $T_{i}$ and $T_{e}$ are not expected to equilibrate. We use the same parameter regime when testing the XMHD algorithm at the ideal MHD limit in this paper.
@@ 3.1 Slope limiter
where $\left(\frac{\lambda_{0}}{\Delta x}\right)^{3}$ is a free parameter and from the shock tube problem, 0.005 is a suitable value. This can be seen akin to a local Knudsen number of around 0.17 which lies in the transition regime away from the continuum limit. Therefore, as the flow density decreases and approaches the molecular limit, the above limiter approaches zero, giving rise to a 1st-order diffusive flux to stabilize the solution by damping out small amplitude but high frequency oscillations. Note that this density-dependent slope limiter is only applicable to problems that contain both high- and low-density regions. For problems in which all densities lie within the low-density regime (e.g., Sect. 4.4), the limiter reduces to 1 after normalization by the reference low density $n_{0}$. Future work could be focused on a more rigorous density-dependent slope limiter without any free parameter.
@@ 3.2 Primitive form for flux Jacobian
Note that the explicit source terms depend on $\mathbf{E}$ and $\mathbf{J}$ which are implicitly updated by the diffusive backward Euler to relax the stiff source term to the equilibrium value where the characteristic speeds of the flux Jacobian become more relevant. Moreover, the relaxation accounts for the speed of light according to Eqs. 51 which shows that larger $\Delta t$ results in faster relaxation since more higher-order terms are damped out. However, these are not sufficient and thus, density-dependent slope limiters are required for the shock tube problem to further damp out small and fast oscillations. More rigorous analysis and testing will be needed in future.
@@ 4.2 Ideal MHD linear wave test
As expected, the results in Fig. 6(a) show that without Hall and resistive terms, the coupling of ideal MHD and other components through the implicit source term update is correct in ensuring that $\mathbf{E}\approx-\mathbf{u}\times\mathbf{B}$ due to the 2nd-order convergence from employing modified PLM and RK2 time integration. However, the entropy wave ceases to exhibit second-order convergence at very fine resolutions. This may be a result of GOL not relaxing to the ideal MHD’s Ohm law as the grid resolution increases. For instance, when different $B_{0}$ and $t_{0}$ are chosen such that $\epsilon_{E}$ is smaller while $\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}$ remains approximately 1 (see Table 2), second-order convergence is observed in Fig. 6(b).
$B_{0}=580$, $t_{0}=10^{-7}$ $B_{0}=58$, $t_{0}=10^{-6}$
$\epsilon_{E}$ $1\times 10^{-9} 1.11×1$$1\times 10^{-11}$
$\epsilon_{J}$ $4.71\times 10^{-10}$ $4.71\times 10^{-10}$
$\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}$ $95\times 10^{-1} 9.9$$95\times 10^{-1}$
Table 2: Different relaxation parameters where the second column is the original dimensional constants used for this problem.
@@
(b) Since entropy wave did not fully converge to 2nd-order especially at high resolution, a different set of dimensional constants are used to ensure that the XMHD solver is able to relax sufficiently to the ideal MHD limit for 2nd-order convergence.
@@ 4.3 Field Diffusion
where $x_{0}=10$cm, $\Delta x=3$cm and Atwood number is set as $A=20$ with a low number density of $n_{0}=10^{18}\text{m}^{-3}$ where Hall effect is significant. This means that the dimensional constants for the PDEs are also adjusted such that $n_{0}=10^{18}$, $L_{0}=1$ and $B_{0}$ set to 0.4568 such that the dimensionless coefficient ($\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}$) is approximately 1. The density profile is then maintained by setting $u_{x}=0$ every timestep. The plasma is also assumed to be isothermal ($T\equivconst$.) and for the ideal MHD components, HLLD Riemann solver [23] is used. Magnetic field is set along the z direction where $B_{0}=0.1\text{T}$ with perturbation of $\delta B_{z}=\delta B\cos\left(\frac{2\pi my}{L_{y}}\right)$ in the y-direction where $\delta B=0.001\text{T}$ and m is the mode number. The plasma beta ($\beta$) is then set to 0.0001 at peak density region where $x=0$. This allows the Hall drift wave to propagate in the y-direction.
@@
Rotor problem is also included to study the onset and propagation of strong torsional Alfven waves in both ideal and non-ideal regimes. Parthenon’s built-in adaptive mesh refinement (AMR), featuring a uniquely designed second-order prolongation and restriction scheme that preserves $\nabla\cdot\mathbf{B}\approx 0$ [36], is also tested with the XMHD solver. This is done using a computational domain of $0.5,0.5]\times[-0.5,0.5]$ with base grid resolution of 100 by 100 and 3 levels of adaptive grid refinement. The heat capacity ratio is set to $1.4$ here. For the ideal MHD case, $\rho_{0}$ is 1 and to test for the increment of non-ideal effects, $\rho_{0}$ is decreased to 0.1. The problem is initialized with
@@ 5 Conclusion
The goal of this work is to develop a semi-implicit finite-volume scheme for solving the XMHD equations while asymptotically preserving the ideal MHD, resistive and Hall limits. This potentially enables the current solver to handle a wider range of density regimes, which is particularly important for high-energy-density (HED) plasma systems. Such systems often exhibit large density variations that must be accurately captured to improve the understanding and control of the plasma’s energy and stability.
@@
This is accomplished by first re-writing the governing equations such that it preserves most of the ideal MHD framework, allowing existing ideal MHD solvers or algorithms to be re-used. For instance, the advantage of preserving locally divergence-free magnetic field by placing magnetic-field at face-center through constrained transport method is employed in this XMHD solver. Naturally, the $\mathbf{E}$ field is placed at cell edges, together with $\mathbf{J}$ field for consistency during implicit source term updates. However, the implicit source term update requires directional splitting to avoid spatial biasing since it requires all of the edge-centered $\mathbf{J}$ and $\mathbf{E}$ components which are not co-located. The Jacobian matrix for the governing equations without the source terms is also obtained to determine the diffusive stabilization for the non-ideal components of the LLF Riemann solver evaluated at the cell faces. But this is not sufficient stability-wise, especially at low density where non-ideal effects become significant and therefore, a density-dependent slope limiter is used to increase flux diffusivity away from the continuum limit. Finally, since all the variables are not co-located, interpolation of values across different topological elements of the computational cell is required and simple arithmetic averaging seems sufficient for the test problems here.
The XMHD solver is then validated using the Brio-Wu shock tube test, producing results consistent with the ideal MHD solver in Athena++. At low densities, the solution departs noticeably from the ideal MHD results but remains stable and free of oscillations. The solver is further tested in the ideal MHD limit by propagating linear MHD waves, which exhibit the expected second-order convergence consistent with the PLM reconstruction and RK2 time integration. Second-order convergence is also confirmed in the resistive MHD limit through the diffusion of a Gaussian magnetic-field profile. In the low density regime corresponding to the Hall MHD limit, the XMHD solver reproduces Hall drift waves in regions with density gradients and captures the linear dispersion relation in agreement with analytical predictions. Lastly, the solver is fully compatible with Parthenon’s AMR framework, preserving the divergence-free condition of the magnetic field via specialized prolongation and restriction operators during refinement.
@@
There are other aspects of the current algorithm that require further understanding and improvement. For example, the solver includes two stiff terms (i.e., $\epsilon_{E}$ and $\epsilon_{J}$) that depend on the dimensional constants used to non-dimensionalize the XMHD equations. These constants must be chosen such that the term $\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}\approx 1$, as this allows one to retain the existing ideal MHD solver without explicitly including the term. The ratio of the timestep $dt$ to $\epsilon_{E}$ and $\epsilon_{J}$ may influence how effectively the system relaxes toward the appropriate physical limit, and future work will involve better understanding this. Furthermore, there are several potential avenues for enhancing the current algorithm such as